OpenStudy (anonymous):

14 to the 7x power = 21

OpenStudy (anonymous):

do you mean \[\large 14^{7x} = 21\]

OpenStudy (anonymous):

YES thanks :D

OpenStudy (anonymous):

turn it into logarithmic form..that will give you \[\Large \log_{14} 21 = 7x\]

OpenStudy (anonymous):

get the value of \(\log_{14} 21\) first then divide by 7...got it?

OpenStudy (anonymous):

yes! So, like Log(14)21 = 24.06... then /7 = 3.43...

OpenStudy (anonymous):

let me check...wait..

OpenStudy (anonymous):

did you do \[\large \log 21 \div \log 14\] on your calculator?

OpenStudy (anonymous):

no i did Log(14)21... that wrong?

OpenStudy (anonymous):

you need to do it like i said...that's how you do logs on calculator \[\large \log_a b = \log b \div \log a\] got it?

OpenStudy (anonymous):

How do I do Log_a

OpenStudy (anonymous):

on a calculator?